OpenStudy (goldrush18):

Suppose Sandra has $16.00 to spend only on ripe bananas and plums. Ripe bananas cost $0.80 each and plums cost $0.20 each. a. write down the algebraic equation for sandra's budget constraint b. if sandra buys only ripe bananas, calculate how many bananas can be bought? c. if sandra buys only plums, calculate how many plums can be bought? d if sandra were to buy 12 ripe bananas, how many plums could be bought with the funds left over? e. if sandra consumes one less ripe banana (i.e 11) how many more plums could be bought? f. is the rate of trade off the same no matter how many ripe bananas are relinquished? g. calculate the opportunity cost of ripe bananas for sandra? h. calculate the opportunity cost of ripe plums for sandra? i. graph the sandra's budget showing the points mentioned in a-d

OpenStudy (anonymous):

are you solving any assignment ?

OpenStudy (goldrush18):

what do you mean by that?

OpenStudy (anonymous):

Suppose: banana price = p1 = 0.8 plum price = p2 = 0.2 banana quantity = q1 plum quantity = q2 A- budget constrain p1.x1 + p2.x2 = 16 B- q1 = 16/p1 = 16/0.8 = 20 C- q2 = 16/p2 = 16/0.2 = 80 D- q1 = 12 Money Spent p1*q1 = 12.0.8 = 9.6 Funds Left: 16 - 9.6 = 6.4 q2 = 6.4/p2 = 6.4/0.2 = 32 E- q1=11 Money Spent p1*q1 = 11.0.8 = 8.8 Funds Left: 16 - 8.8 = 7.2 q2 = 7.2/p2 = 7.2/0.2 = 36 More 4 plums F- I don't know this term, relinquished… I G- Opp Cost of Bananas = How many plums she has to relinquish to buy one more plum? Using all her budget, she can have: Only buys bananas: q1= 20, q2=0 Only buys plums : q1=0, q2=80 Opp Cost of bananas in terms of plums: 0-20/80-0= - 0.25 She has to abdicate of 0.25 bananas to buy one more plum. H- Opp Cost of Plums = How many bananas she has to relinquish to buy one more banana? Using all her budget, she can have: Only buys bananas: q1= 20, q2=0 Only buys plums : q1=0, q2=80 Opp Cost of plums in terms of bananas: 0-80/20-0= - 4 She has to abdicate of 4 plums to buy one more banana. I- Easy

OpenStudy (anonymous):

Ops A- budget constrain p1.q1 + p2.q2 = 16

OpenStudy (anonymous):

F - I think relinquish is to give up right? Well. if the amount of bananas you have to give up to get more plums change, the rate of trade-off changes as well.. SOrry if i told something wrong.. Bye

OpenStudy (anonymous):

Klanfer - The rate of trade off remains the same at 0.25 bananas/plum, or 4 plums/banana. The costs do not change with the number of bananas bought.

OpenStudy (anonymous):

Sure I agrre with you Mysthalin. HOwever, note that I did not say that the cost will change as she buys more or less bananas... What I said is that the cost will change, if the amount of bananas she needs to give up to buy more plums changes... In other words, if the numbers of bananas she needs to give up to buy more plums changes, the trade-off rate will be different.. since this is the definition of trade-off... See you!

OpenStudy (anonymous):

If neither costs change, that means that the trade-off rate will remain the same, as the number of plums/bananas you need to give up will also remain the same.